# Vision-Language Models can Identify Distracted Driver Behavior from Naturalistic Videos

### 04 Jan 2024 v3

@@ I Introduction
DIstracted driving accounts for almost 10% of all fatalities on U.S. roads. Although cell phone usage is often synonymous with distracted driving, the problem is broader. The Center for Disease Control and Prevention (CDC) identifies three forms of distraction: i) visual distraction, where the driver takes their eyes away from the road. This can be due to several reasons, such as looking at the GPS, Billboard, etc.; ii) manual distraction, where the driver takes their hands off of the steering wheel; and iii) cognitive distraction, where the driver takes their mind off the driving task, which can be due to singing, talking etc. The association of crashes with distracted behavior is underreported. Verifying the distraction in the post-crash investigation is difficult, and the driver at fault might not disclose this information voluntarily. Contrary to the numbers reported in crash analysis, Naturalistic Driving Data analysis by Dingus et al. [1] reported that almost 90% of their studied 905 crashes, which resulted in injury or property damage, can be attributed to driver-related factors (i.e., error, impairment, fatigue, and distraction). Naturalistic driving study (NDS) research involves observing drivers in real-world settings using unobtrusive instruments fitted to their vehicles. These studies cover drivers’ daily driving activity for a few months to years. With low-cost sensors and robust data handling infrastructure, NDS research is now easier to implement and has paved the way to comprehensively understanding the occurrence of driving events and driver responses in a natural environment. In most recent studies, such analysis has required humans to annotate driver distraction manually. The manual approach is useful but very time-consuming and significantly increases the analysis time. This study aims to reliably and efficiently identify distracted driving actions by drivers from naturalistic driving videos.
With the significant advances in deep learning and computer vision, there has been a considerable surge in video analytics for the classification of distracted driving. In recent studies, convolutional neural networks (CNN) have been widely employed for video-based distracted driving detection. For instance, Moslemi et al. [2] proposed a 3D CNN model for distracted driving detection using a dataset composed of videos captured in a naturalistic driving study. This model was designed to learn spatiotemporal features from video sequences, thus allowing it to capture not only the appearance of distracted behaviors but also their temporal dynamics. However, this type of model often requires large amounts of labeled training data, which is both time-consuming and costly to collect and annotate. Therefore, researchers have recently begun to investigate unsupervised [3] and semi-supervised [4] learning techniques. For instance, Xie et al. [5] introduced a semi-supervised learning method for distracted driving detection, wherein they combined labeled and unlabeled data to train their deep learning model.
The recent development of vision-language multimodal pretraining frameworks showcases promising potential across diverse applications [6, 7, 8, 9]. These pretraining frameworks employ a multi-modal contrastive-loss (see CLIP [10]), which paves the way for creating robust embeddings encompassing both text and vision tasks concurrently. This approach allowed several domains to use an almost zero-shot approach for tasks, such as object classification and object detection from images [10, 11, 12, 13]. The study presented in this paper leverages CLIP’s vision and language embeddings for the specific task of distracted driving action recognition.
@@
Distracted driving action recognition involves analyzing video data captured from a camera positioned inside the car, which records extensive time frames of driver behavior. The goal is to utilize this video data to determine whether the driver is distracted. Intuitively, one might consider training a video-text model from the ground up for this task. However, directly training such a model is impractical due to the need for large-scale video-text training data and considerable GPU resources. A more viable approach is to adapt pre-existing language-image models to the video domain [14, 15, 16]. Several recent studies have investigated strategies for transferring knowledge from these pretrained language-image models to other downstream tasks [12, 17, 18, 19, 20].
Our proposed method benefits from key aspects of the pretrained CLIP model [10] and provides two main advantages. The first advantage is its ability to utilize zero-shot transfer, obviating the need for intensive training. This feature enhances scalability and feasibility, allowing for a natural language-supervised understanding of video activities rather than relying exclusively on traditional supervised approaches. Unlike conventional computer vision methods, which tend to memorize human-annotated labels or features often found non-transferable, this model generally exhibits superior performance across diverse datasets rather than constrained to a specific dataset. The second advantage of the proposed method is the potential to further finetune the visual representation with minimal annotations and training, thereby enhancing the system’s overall performance. In addition, it allows us to assess the generalizability and robustness of the finetuned models in diverse real-world settings.
@@
    We introduce and explore the performance of various vision-language-based frameworks (adapted from [10]) for distracted driving action recognition tasks with minimum training. Moreover, we propose a unique validation method named ‘leaving a set of driver out’, which involves subject-wise separation during training to prevent overfitting to specific driver characteristics. It also tests the model’s generalizability to previously unseen driver behaviors.
@@
    We present both frame-based and video-based models and show that including temporal information in the existing vision-language models offers superior performance on several distracted driving datasets. Moreover, we execute an exhaustive fine-tuning with extensive hyperparameter tuning across multiple datasets to achieve the optimized performance on our proposed models. This confirms the models’ effectiveness in recognizing driver distractions under a variety of real-world conditions.
@@
    We perform an in-depth analysis of the models’ performance across different datasets and various experimental settings. We demonstrate the model’s robustness and strong generalization ability by maintaining high-performance levels even with reduced training data sizes. These attributes underscore the potential for deploying our models in real-world scenarios where training data may be limited.
In this study, we have developed and tested four unique frameworks on four distinct datasets that are tailored specifically for analyzing distracted driving behavior. These frameworks include both frame-based and video-based models. The rest of this paper is organized as follows: Section II provides a review of relevant research on distracted driving action recognition, highlighting key contributions as well as gaps that our study aims to address. Section III introduces the details of our methods. Section IV provides details about the datasets and the experimental setup for each framework. Section V presents the results of the performance evaluation of the proposed frameworks, performance comparison with existing models for distracted driving, an investigation into the impact of different real-world settings on the model performance and an ablation study. The paper concludes in Section VI.
@@ II Related Work
In recent years, the area of driver action recognition and behavior analysis has attracted significant research interest. Various methods have been proposed including -
II-A Machine Learning-based Techniques
Prior to recent advances in computer vision, traditional machine learning approaches primarily relied on manually selecting key features and pairing them with a classifier [21]. Features were collected from the driver’s hand, head-pose, face, body-pose, eye gaze [22, 23, 24, 25, 26, 27] or sensors [28, 29]. The choice for common classifier includes SVM [30], HMM [31], random forest [32], KNN [33] and so on. However, the major limitation of these methods is that the manual feature selection lacks generalization, making it challenging to apply them effectively across diverse driving scenarios [34]. Many of these techniques are specifically tailored to certain datasets, which lack broad applicability due to the lack of end-to-end training on large-scale data.
II-B Supervised Learning-based Techniques
@@
The emergence of camera sensors and advancements in Convolutional Neural Networks (CNNs) have significantly influenced the field of computer vision and driver observation. As a result, a significant portion of the current methodologies depended on various supervised CNN models [35, 36, 37] for extracting important features from image and video content. Some CNN-based frameworks gather features from facial recognition [38], optical flow [39], and body posture assessment [40] for distraction action recognition. Some of the most common CNN backbones used in literature are VGG16 [41], ResNet50 [42], AlexNet [43], Inception-v3 [44], MobileNet-v2 [45], ShuffleNet-v2 [46] etc. Various CNN approaches are employed based on the availability of dimensions (2D/3D), modalities (RGB, IR, Depth, [36, 35]), camera views (front, side etc) of datasets and ensemble techniques [47]. Traditional video-based action recognition CNN models often evolve from image-based 2D CNN models, adapting to incorporate the temporal aspect of videos. Strategies for integrating the temporal dimension include using 2D CNNs on multiple frames or 3D CNNs on video clips [48, 49]. However, these supervised CNN-based approaches require vast amounts of labeled data for training, making the process resource-intensive. This heavy reliance on extensive labeled data poses a significant challenge and can limit their generalization capabilities [3]. Furthermore, while 3D CNNs can learn temporal features from video sequences, their implementation is often limited by high computational requirements and real-world challenges in processing naturalistic videos [50, 51, 48]. Another technique that is commonly adopted involves ensembling multiple CNN networks [52, 53, 54]. Here, each CNN works as a unique feature extractor. Finally, they are incorporated into a weighted ensemble to yield a comprehensive global feature. In addition, there are methods that adopt feature-level fusion [48, 55, 56]. These methods separately learn from RGB appearance and time-based features and later combine the learned features.
II-C Weakly-Supervised Learning-based Techniques
@@
The landscape of distracted action recognition is increasingly being shaped by Vision Transformer (ViT)-based techniques [57, 58, 59, 60], which have started to dominate in video action recognition recently. A ViT-based, weakly supervised contrastive (W-SupCon) learning approach is proposed in [4], where it quantifies distracted behaviors by measuring their deviation from the normal driving representations. A key limitation of this study is the challenge of distinguishing high-risk driving behaviors from normal ones due to their similar physical movements indicating limited generalization.
II-D Unsupervised Learning-based Techniques
A multilayer perceptron-based unsupervised model is proposed in [3] where a ResNet50 [42] backboe is trained for enhanced feature extraction. However, this approach’s generalizability is limited, having been tested on only one dataset without large-scale validation.
@@ II-D Unsupervised Learning-based Techniques
The recent progress in vision-language pretraining approaches [19, 12, 61] has demonstrated improved performance in tasks related to video action recognition tasks. These models are developed on a widely used vision-language model named CLIP [10]. For video-based CLIP training, Miech et al. [62] introduce a large dataset referred to as HowTo100M. Additionally, for learning the visual-semantic joint embedding, Mithun et al. [63] propose a two-stage approach that leverages web images and corresponding tags. Dong et al.  [64] attempt to solve the challenging problem of zero example video retrieval using textual features. Therefore, different methods have been developed to effectively learn from and utilize visual and textual data for tasks and visual-semantic joint embedding, which underlines the potential of integrating visual and textual data in developing advanced models. While the CNN and ViT-based models exhibit satisfactory results on specific datasets, their effectiveness is significantly tied to large-scale data and substantial efforts for data annotation. Additionally, the training of these models is often geared towards mapping predefined categories, which impedes their ability to learn adaptable visual features on new data. Therefore, rather than confining the learning to specific visual features within a single modality framework, learning from natural language supervision could provide increased adaptability and wider applicability [10]. Our method leverages the pretrained CLIP model to offer two significant benefits. Firstly, it enables zero-shot transfer, eliminating the requirement for extensive dataset-specific training. Secondly, it allows further finetuning of visual representations with only a small amount of training data. Moreover, our approach addresses the overfitting issue encountered in previous studies where subject-level separation in training and testing sets was not maintained. We discuss this step in detail in section III
@@ III Methodology
Figure 1: An overview of the image-based and video-based CLIP frameworks. The yellow box shows the sampled frames from a 10-second event, i.e., a total of 300 frames at a rate of 30 frames per second. These frames are sampled from raw data (as depicted in Fig. 3). The VideoCLIP model uses 30 consecutive frames to form a short video clip as input, while the Single-frameCLIP model takes a single frame at a time. For the first sampled frame, the VideoCLIP generates a 30-frame video clip by repeating Image-1 sixteen times and including Images 2 to 15 (a total of 30 frames in the green box). The SingleframeCLIP model only uses Image-1 as its first input. This process continues for each subsequent frame. Finally, predictions for an entire event are made by calculating the majority vote from all individual frame predictions. Note that the Multi-frameCLIP applies majority voting on SingleframeCLIP’s predictions to get the final event prediction.
@@
The primary objective of this work is to understand distracted driving behaviors captured from in-vehicle video footage and build an effective driver monitoring system. While distractions during driving can span a variety of situations, we predominantly focus on actions that disrupt the driver’s attention from the primary task of safely operating the vehicle on the road. In addition, we only focus on a single distracted action that occurred for a certain duration and consider the action as a distracted event. We also assume that there are no multiple distracted events occurring simultaneously. Examples of distracted events are talking on the phone, reaching behind, or drinking water while driving. Visual data associated with distracted driving can present several complications in detecting distracted driving, including low-quality video data and the extensive human effort required to annotate videos for training supervised models. Furthermore, without a large volume of annotated visual data, training a high-performance supervised network is impossible. Therefore, we investigate to create a generalized framework that performs well without a strict reliance on labeled data or with only limited labeled data. To achieve this, we propose to leverage a contrastive vision-language pretraining framework designed to tackle these challenges with minimal annotated data. Moreover, we pay special attention to how our experiments were conducted, ensuring a robust and generalizable testing strategy. In real-world scenarios, the model will be expected to generalize its learning to new and unseen drivers. Therefore, we employed two combined methods: 1) subject-wise separation - ‘leaving a set of driver out’ method where we separate out a unique set of driver samples only for testing. 2) subject-wise cross-validation - $k$-fold cross-validation method where we implement $k$-fold cross-validation technique individually to each driver set. For the first approach, the test split drivers’ data are never presented to the model during its training phase. It ensures that the model can make inferences fully on unseen data during testing. These two methods ensure zero overlapping between training and testing splits and zero memorizing across all the runs we conducted. The average accuracy obtained from all the $k$-fold results is reported as the final accuracy of our frameworks. This process allows us to accurately assess the overall performance of our proposed distracted driving models across all the datasets.
@@
Furthermore, we recognize the significance of addressing multiple distracted driver behaviors occurring simultaneously or alternately. Our observations have revealed that the effectiveness of handling such scenarios is dependent on the specific datasets used and the formulation of prompts. It’s essential to underscore that dealing with multiple distracted behaviors is indeed a complex challenge within the scope of this problem. While our current framework has not been specifically tested on scenarios involving multiple distracted behaviors, we firmly believe that, with the availability of appropriate datasets and prompt customization, our framework can be adapted to successfully detect and analyze such situations. We acknowledge that this aspect of the problem requires further research and exploration.

### 21 Mar 2024 v4

@@ I Introduction
DIstracted driving accounts for 8% of fatal crashes, 14% of injury crashes, and 13% of all police-reported motor vehicle traffic crashes in 2021 on U.S. roads. It has emerged as a significant public safety concern and contributed to a substantial number of road accidents worldwide [1]. The World Health Organization reported that 1.19 million people died in traffic accidents worldwide in 2023 [2]. Although cell phone usage is often synonymous with distracted driving, the problem is broader. The Center for Disease Control and Prevention (CDC) identifies three forms of distraction [3]: i) visual distraction, where the driver takes their eyes away from the road. This can be due to several reasons, such as looking at the GPS, Billboard, etc.; ii) manual distraction, where the driver takes their hands off of the steering wheel; and iii) cognitive distraction, where the driver takes their mind off the driving task, which can be due to singing, talking etc. The association of crashes with distracted behavior is underreported [4]. Verifying the distraction in the post-crash investigation is difficult, and the driver at fault might not disclose this information voluntarily. Among all the U.S. drivers involved in fatal crashes in 2021, 5% were reported as distracted at the time of the crashes [1]. Contrary to the numbers reported in crash analysis, Naturalistic Driving Data analysis by Dingus et al. [4] reported that almost 90% of their studied crashes, which resulted in injury or property damage, can be attributed to driver-related factors (i.e., error, impairment, fatigue, and distraction). Numerous studies have explored the relation between precrash driver instability and distracted driving [5, 6, 7]. Arvin et al. [6] observed a substantial relation between distracted driving with crash intensity. They showed the probability of a severe crash increased by 11.1% due to driver distractions. Therefore, real-time detection of the driver’s status and providing effective feedback signals to minimize distraction is essential. It could significantly decrease the number of accidents resulting from distracted driving [8, 9]. In recent years, with the rapid growth of affordable sensors and storage technologies, Naturalistic Driving Studies (NDS) have become the most common approach for analyzing driver behavior [10]. NDS offers video data which is preferred for investigating driver distractions in unbiased real-world settings. These studies leverage computer vision techniques to detect distracted driving, aiming to reduce crash risks by alerting drivers autonomously. However, the reliance on human annotation of driver distraction in these studies introduces significant time delays. Our study aims to efficiently identify the common distraction activities in naturalistic driving videos using a vision-language-based fremework [11] with limited annotated data, reducing training time and computational resources.
Recent advancements in deep learning and computer vision have significantly surged video analytics to identify distracted driving. In recent studies, the Convolutional Neural Networks (CNN) have been widely employed [12, 13, 14, 15, 16] for video-based distracted driving detection. However, these supervised CNN-based approaches require vast amounts of labeled data for training, making the process resource-intensive. This heavy reliance on extensive labeled data poses a significant challenge and can limit their generalization capabilities. Therefore, researchers have recently started to investigate unsupervised [17] and semi-supervised [18] learning techniques.
Most of the distracted driving recognition tasks in computer vision primarily rely on analyzing in-vehicle video footage to detect driver inattentiveness [10, 13]. The challenge lies in the model’s ability to handle diverse and unexpected driving behaviors that are common in real-world driving scenarios. Therefore, it is essential for the models to efficiently adapt and recognize a broader spectrum of actions, particularly focusing on those that are rare or challenging to classify. While traditional CNN models perform well on specific datasets, their success depends heavily on large-scale data and annotation efforts [17]. Additionally, training of these models is often geared towards mapping predefined categories, which impedes their adaptability to new data. Therefore, rather than confining the learning to specific visual features within a single modality framework, learning from natural language supervision could provide adaptability and wide applicability [11, 19, 20]. The recent advancements in vision-language pretraining frameworks showcase significant promise for diverse applications [21, 22, 23, 24]. These frameworks utilize a multi-modal contrastive loss to generate robust embeddings that simultaneously address text and vision tasks. This approach allows several domains to use an almost zero-shot strategy for image-based object classification and detection tasks [11, 25, 20, 26].
@@
The study presented in this paper leverages CLIP’s [11] vision and language embeddings for identifying distracted driving actions. However, directly training such a model for video analysis is impractical due to the need for large-scale video-text training data and considerable GPU resources. An alternative and more feasible strategy is leveraging existing language-image models for application in the video domain [27, 28, 29]. Several recent studies have investigated strategies for transferring knowledge from these pretrained language-image models to other downstream tasks [30, 31, 32].
Our proposed method leverages the pretrained CLIP encoders and offers two significant benefits. Firstly, the zero-shot transfer eliminates the requirement for extensive dataset-specific training, making it more scalable and practical for understanding video content through natural language. This contrasts with traditional methods that rely on memorizing human-annotated labels or features which are often non-transferable. Secondly, it allows further finetuning of visual representations with a small amount of training data, thus enhancing the model’s performance. Moreover, our approach addresses the evaluation issue encountered in previous studies [33] by introducing cross-validation techniques and subject-level separation in training and testing sets.
@@
    We introduce and explore the performance of various vision-language-based frameworks (adapted from [11]) for distracted driving action recognition tasks with minimum training. We present both frame-based and video-based models and show that including temporal information in the existing vision-language models offers superior performance on several distracted driving datasets. Moreover, we execute exhaustive fine-tuning with extensive hyperparameter tuning across multiple datasets to achieve the optimized performance. It tests the model’s generalizability to previously unseen drivers.
@@
    We propose a unique validation method named ‘leaving a set of driver out’, which involves subject-level separation during training to prevent overfitting to specific driver appearance. Moreover, we perform a cross-validation step to confirm the models’ effectiveness in a variety of real-world conditions.
@@
    We perform an in-depth analysis of the models’ performance across different datasets and various experimental settings. Additionally, we demonstrate the model’s robustness by maintaining high-performance levels even with reduced training data sizes. These attributes underscore the potential for deploying our models in real-world scenarios where training data may be limited.
In this study, we develop and test four unique CLIP-based frameworks on four distinct distracted driving datasets. These frameworks include both frame-based and video-based models.
The rest of this paper is organized as follows: Section II provides a review of relevant research on distracted driving action recognition. Section III includes the details of our methods. Section IV provides details about the datasets and the experimental setup. Our results and observations are presented in Section V. The paper concludes in Section VI.
@@ II Related Work
In recent years, the area of driver action recognition and behavior analysis has attracted significant research interest. Generally, we notice two distinct categories of research works in the literature.
II-A Visual Foundation methods
@@
Before the advent of computer vision techniques, distracted driving detection primarily relied on traditional machine learning methods, which involved manually selecting key features [34, 35, 36, 37, 38] and pairing them with a simple classifier [39, 40, 41, 42]. However, the major limitation of these methods is that the manual feature selection lacks generalization, making it challenging to apply them effectively across diverse driving scenarios [10, 43]. However, the emergence of camera sensors and advancements in Convolutional Neural Networks (CNNs) have significantly influenced the field of driver behavior observation. Therefore, a significant portion of the current methodologies depends on various supervised CNN models [12, 13, 14] and Deep Neural Networks (DNN) [44, 43] for extracting important features from the image and video content. Some CNN-based frameworks gather features from facial recognition [45], optical flow [46], and body posture assessment [47] for distraction action recognition. Some of the most common CNN backbones used in literature are VGG16 [48], ResNet50 [49], AlexNet [50], Inception-v3 [51], MobileNet-v2 [52], ShuffleNet-v2 [53] etc. However, the parameter size of CNN poses a challenge for practical applications [43]. Therefore, some lightweight CNN frameworks [44, 54] are proposed in the literature. Various CNN approaches are employed based on the availability of dimensions (2D/3D), modalities (RGB, IR, Depth, [13, 12]), camera views (front, side etc.) of datasets and ensemble techniques [55]. Traditional video-based action recognition CNN models often evolve from image-based 2D CNN models, adapting to incorporate the temporal aspect of videos. Strategies for integrating the temporal dimension include using 2D CNNs on multiple frames or 3D CNNs on video clips [56]. While 3D CNNs can learn temporal features from video sequences, their implementation is often limited by high computational requirements and real-world challenges in processing naturalistic videos [57, 58, 56]. Another technique that is commonly adopted involves ensembling multiple CNN networks [59, 60, 61]. Here, each CNN works as a unique feature extractor. Finally, they are incorporated into a weighted ensemble to yield a comprehensive global feature. In addition, there are methods that adopt feature-level fusion [56, 62, 63]. These methods separately learn from RGB appearance and time-based features and later combine the learned features.
II-B Distracted Driving Recognition methods
@@
Before the widespread use of Vision Transformer (ViT) models [64], some works relied on CNN for distracted driving detection. For instance, Moslemi et al. [16] proposed a 3D CNN model, which was pretrained on both the ImageNet and Kinetics datasets and fine-tuned on a distracted driving detection dataset. This model was designed to learn spatiotemporal features from video sequences. Recently, the landscape of distracted action recognition is increasingly being shaped by ViT-based techniques [65, 66, 67, 68], which have started to dominate video action recognition tasks. A ViT-based, weakly supervised contrastive (W-SupCon) learning approach is proposed in [18], where it quantifies distracted behaviors by measuring their deviation from the normal driving representations. A key limitation of this study is the challenge of distinguishing high-risk driving behaviors from normal ones due to their similar physical movements indicating limited generalization. A multilayer perceptron-based unsupervised model is proposed in [17] where a ResNet50 [49] backbone is trained for enhanced feature extraction. However, this approach’s generalizability is limited, having been tested on only one dataset without large-scale validation. Moreover, a recent study by Chai et al. [33] points out that many approaches in distracted driving detection research do not undergo thorough performance evaluation, leading to potentially inflated accuracy claims [10, 69, 70]. They highlight the significance of implementing driver-level separation and cross-validation techniques during evaluation to address these shortcomings.
@@
The recent progress in vision-language pretraining approaches [19, 20, 71] has demonstrated improved performance in tasks related to video action recognition. These models are developed on a widely used vision-language model named CLIP [11]. For video-based CLIP training, Miech et al. [72] introduce a large dataset referred to as HowTo100M. Additionally, for learning the visual-semantic joint embedding, Mithun et al. [73] propose a two-stage approach that leverages web images and corresponding tags. Dong et al.  [74] attempt to solve the challenging problem of zero example video retrieval using textual features. Therefore, various methods have been developed to effectively learn from the textual and visual-semantic joint embeddings. These methods underline the potential of integrating visual and textual data in developing advanced multimodal models.
@@ III Methodology
Figure 1: An overview of the image-based and video-based CLIP frameworks. The yellow box shows the sampled frames from a 10-second event, i.e., a total of 300 frames at a rate of 30 frames per second. These frames are sampled from a raw video. The VideoCLIP uses 30 consecutive frames to create a short video clip as input, while the Single-frameCLIP takes a single frame at a time. For the very first sample, the VideoCLIP generates a video clip by repeating Image-1 sixteen times and including Images 2 to 15 (a total of 30 frames in the green box). The SingleframeCLIP only uses Image-1 as its first input and makes the prediction. This process continues for each subsequent frame. Note that the prediction for an entire event is made by calculating the majority vote from all individual frame predictions. The Multi-frameCLIP applies majority voting on SingleframeCLIP’s predictions and VideoCLIP applies majority voting on its output to get the final event prediction.
@@
This study aims to understand and analyze distracted driving behaviors from in-vehicle video footage and build an effective driver monitoring system. Despite the wide range of driving distractions, we only focus on specific actions or events that divert the driver’s attention from safe driving. Furthermore, we consider each unique distracted action that lasts for a certain duration as an individual event. We also assume that these distracting events do not overlap or occur simultaneously. For example, distracting events can be talking on the phone, reaching behind, or drinking water while driving. Detecting these distracting events from video data involves several challenges, including low-quality video and the extensive human effort required to annotate videos or images for training supervised models. Therefore, we aim to develop a generalized framework that performs well without strictly relying on labeled data or with only limited labeled data. Our proposed framework is based on a contrastive vision-language pretraining approach designed to address these challenges with minimal annotated data. Moreover, we carefully conduct our evaluation experiments to ensure a robust and generalizable testing strategy. In real-world scenarios, the model will be expected to encounter new and unseen distracting events. Therefore, we employ two methods during training: 1) subject-wise separation method - ‘leaving a set of driver out’ where we separate out a unique set of driver samples only for testing. 2) subject-wise cross-validation method - $k$-fold cross-validation where we implement $k$-fold cross-validation technique individually to each driver set. For the first method, the “test split drivers’ data” are never presented to the model during its training phase. It ensures that the model only focuses on distracting events rather than the subject’s appearance and can make predictions fully on unseen data during testing. These two methods ensure zero overlapping between training and testing splits and zero memorizing across all the runs we conducted. The average accuracy obtained from all the $k$-fold results is reported as the final accuracy of our frameworks. This process allows us to accurately assess the overall performance of our proposed models across all the datasets.
@@
Furthermore, we recognize the significance of addressing multiple distracting events occurring simultaneously or alternately. Our observations have revealed that the effectiveness of handling such scenarios depends on the quality of the datasets and the formulation of the prompts. It is essential to underscore that dealing with multiple distracted behaviors is indeed a complex challenge within the scope of this problem. While our current framework has not been specifically tested on multiple distracting events, we firmly believe that, with the availability of appropriate datasets and prompt customization, our frameworks can be adapted to detect and analyze such situations. We acknowledge that this aspect of the problem requires further research and exploration.